OpenStudy (anonymous):

solve. then round answer to the nearest hundredth. 10^x+4=1000

OpenStudy (anonymous):

@dpflan

OpenStudy (anonymous):

Is it\[10^{x+4} = 1000\] In anyway case, it's just like the last one, you can solve this one in a minute ;)

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

10^x+4=1000 10^x = 996 x = log 996 2.9982

OpenStudy (anonymous):

wait, if it's \[10^{x+4}\] then you need to re-solve

OpenStudy (anonymous):

Otherwise, you got it

OpenStudy (anonymous):

so if it is 10^x+4 its wrong?

OpenStudy (anonymous):

No, you solved it correctly if the +4 is not in the exponent

OpenStudy (anonymous):

be careful of what you're solving... which of these is the correct expression on the left: \[\large 10^{x+4} \] or \[\large 10^{x}+4 \]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

then like @dpflan , said, you need to resolve... your original solution is for the second expression

OpenStudy (anonymous):

10^(x + 4) = 1000 x + 4 = 3 x = -1

OpenStudy (anonymous):

yeah, nice!

OpenStudy (anonymous):

how would i round that to the nearest hundredth though?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Well, it currently have no fractional value, nothing passed the decimal point, so you can rewrite that \[-1=-1.00000....\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem, thanks for the medal!